OpenStudy (otonogold):

I do not know how to simplify this expression -5+i / 2i

OpenStudy (anonymous):

is it... \[-5 + \frac{ i }{ 2i }\]

OpenStudy (anonymous):

@OtonoGold ?

OpenStudy (anonymous):

ok i assume it is correct... we put expression in fraction by making 2i as common denominator for both terms... \[\frac{ -10i +i }{ 2i }=\frac{ -9i }{ 2i }=-4.5\]